OpenStudy (naveenbhatia1312):

Find all solutions in the interval [0, 2π). (6 points) sin^2x - cos^2x = 0

OpenStudy (danjs):

what you try already

OpenStudy (naveenbhatia1312):

confused

OpenStudy (danjs):

try using one of the identity from the last prob we did, replace cos^2 with it and solve for what sin will be

OpenStudy (danjs):

\[\sin^2(\theta)- [1 - \sin^2(\theta)] = 0\] solve that for sin(theta)

OpenStudy (anonymous):

can I help

OpenStudy (anonymous):

any one

OpenStudy (naveenbhatia1312):

yes please

OpenStudy (naveenbhatia1312):

@damirax

OpenStudy (anonymous):

I a here

OpenStudy (naveenbhatia1312):

I need help with the problem above

OpenStudy (anonymous):

sorry my comp glitching

OpenStudy (anonymous):

is this flvs

OpenStudy (naveenbhatia1312):

no

OpenStudy (anonymous):

then why are there points (6 points) I cant help you if it is not from a book soooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooory

OpenStudy (anonymous):

here is a hint next time delet points